OpenStudy (anonymous):

If (x) =-2x-4 Calculate the following A) f(-1) C) f(2) +g(-4)

OpenStudy (anonymous):

There's no equation for g(x)

OpenStudy (anonymous):

-2x-4 f(1), let x=1 -2(1)-4= ?

OpenStudy (anonymous):

G(5)

OpenStudy (anonymous):

what is the equation for g of x?

OpenStudy (anonymous):

G(x) =-2x-4